OpenStudy (anonymous):

Really confused on this. Please help! What is the pH of a solution whose [H3O+] is 1 X 10 -5 M? What is the pH of a solution whose [H3O+] concentration is 3 X 10-3 M? What is the pH of a solution with a [H3O+] concentration of 1 X 10 -12 M?

OpenStudy (anonymous):

\[\color{green}{\text{pH} = -\log [H_3O^{+}]\\ = - \log(1 \times 10^{-5})}\\\text{etc} \ldots\]